### v2

# RankCLIP: Ranking-Consistent Language-Image Pretraining

Abstract
Self-supervised contrastive learning models, such as CLIP, have set new benchmarks for vision-language models in many downstream tasks. However, their dependency on rigid one-to-one mappings overlooks the complex and often multifaceted relationships between and within texts and images. To this end, we introduce RankCLIP, a novel pretraining method that extends beyond the rigid one-to-one matching framework of CLIP and its variants. By extending the traditional pair-wise loss to list-wise, and leveraging both in-modal and cross-modal ranking consistency, RankCLIP improves the alignment process, enabling it to capture the nuanced many-to-many relationships between and within each modality. Through comprehensive experiments, we demonstrate the effectiveness of RankCLIP in various downstream tasks, notably achieving significant gains in zero-shot classifications over state-of-the-art methods, underscoring the importance of this enhanced learning process. Code and model checkpoints are available at https://github.com/Jam1ezhang/RankCLIP.
RankCLIP: Ranking-Consistent Language-Image Pretraining
Yiming Zhang††thanks: These authors contribute equally. Correspondence to Zenghui Ding <dingzenghui@iim.ac.cn>. Zhuokai Zhao Zhaorun Chen2 Zhili Feng3 Zenghui Ding1 Yining Sun1 1HFIPS, Chinese Academy of Sciences 2University of Chicago 3Carnegie Mellon University 4University of Science and Technology of China
1 Introduction
In the realm of computer vision (CV) (Voulodimos et al., 2018), natural language processing (NLP) (Chowdhary and Chowdhary, 2020), and multimodal deep learning (Jabeen et al., 2023; Zhao et al., 2023), the alignment between visual and textual modalities (Singh et al., 2022; Chen et al., 2024a) has emerged as a cornerstone for downstream applications, ranging from image captioning (Ghandi et al., 2023) to zero-shot classification (Pourpanah et al., 2022). Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021) marks a significant advancement in this field, demonstrating incredible performance from training on large amounts of text-image pairs to create self-supervised models that understand (Hendrycks et al., 2021a, b; Chen et al., 2024b) and generate (Ramesh et al., 2021; Crowson et al., 2022) descriptions of visual contents.
Following the success of this contrastive learning paradigm, many recent works have been developed upon the original CLIP. More specifically, these enhancements focus on optimizing data efficiency through intrinsic supervision (Li et al., 2021), as well as improving downstream performance via cross-modal late interaction (Yao et al., 2021), hierarchical feature alignment (Gao et al., 2022), geometric consistency regularization (Goel et al., 2022), additional learning (Mu et al., 2022), adaptive loss (Yang et al., 2023), hierarchy-aware attentions (Geng et al., 2023), and softer cross-modal alignment (Gao et al., 2024).
Despite the improvements, these methods often have reliance on strict pairwise, cross-modal, and one-to-one mappings between images and texts, overlooking the actual many-to-many relationships that exist both cross-modal and in-modal in real-world data (Chun, 2023). For example, as shown in Fig. 1, while pretrained models like CLIP can correctly classify dog, cat and airplane, they do not necessarily learn that dog and cat are more close to each other than dog and airplane, in terms of both in-modal (dog text is more similar to cat text than to airplane text) and cross-modal (dog text is more matched to cat image than to airplane image) similarities. Because it is rooted from the current contrastive loss that only the correct pairs are optimized while the rest of the unmatched pairs are treated the same, resulting in a large amount of information not used and unknown to the model during and after the training process.
Recognizing the complex many-to-many relationships as well as the rich information contained within both in-modal and cross-modal data, we introduce Ranking-Consistent Language-Image Pretraining, (RankCLIP), which employs ranking consistency to learn and optimize similarity levels both between (cross-modal) and within (in-modal) the text-image pairs. The concept of ranking consistency stems from the simple observations that similar texts often correlate with similar images, as seen with the dog, cat and airplane example in Fig. 1. It effectively captures secondary similarity relationships among unmatched pairs, enabling the model to learn more efficiently for free compared to relying solely on matched pairs. Ranking consistency is conveniently modeled as an additional loss term to the traditional contrastive loss, requiring no extra external modules. It acts as a plug-and-play improvement for many existing methods, including those focusing on data-efficiency (Li et al., 2021), potentially boosting performance in both efficiency and effectiveness.
The main contributions of this paper are: (1) RankCLIP, a novel contrastive pretraining method that uses ranking consistency to exploit the many-to-many relationships within data, thereby enhancing performance in downstream tasks such as zero-shot classification and retrieval accuracy; and (2) through comprehensive experiments conducted on multiple datasets, we demonstrate the superior effectiveness of RankCLIP in improving pretraining model performance without requiring any additional data or computational resources.
2 Related Work
Vision-language pretraining has witnessed significant advancements over the past years (Chen et al., 2023; Du et al., 2022; Long et al., 2022). Models such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021) and FLAVA (Singh et al., 2022) have pioneered the contrastive learning paradigm applied with text-image pairs, showcasing remarkable performance and robustness in downstream tasks. Many follow-up works, mostly built upon CLIP, have been proposed since then. Li et al. (2021) introduced DeCLIP, improving zero-shot performance through intrinsic supervision. FILIP (Yao et al., 2021) advances CLIP’s alignment between image patches and text with a cross-modal interaction mechanism. Gao et al. (2022) developed PyramidCLIP, using hierarchical feature alignment to boost model efficiency and performance. Additionally, SLIP (Mu et al., 2022) merges self-supervised learning with CLIP pre-training for improved visual representation and accuracy. Goel et al. (2022) introduced CyCLIP, augmenting CLIP with geometric consistency regularizers to enhance robustness and performance under varied conditions.
Recently, Yang et al. (2023) introduced ALIP, an adaptive pre-training model that enhances language-image alignment using raw text and synthetic captions with dynamic adjustments. HiCLIP (Geng et al., 2023) refines CLIP by adding hierarchy-aware attentions to uncover semantic hierarchies in images and texts. EqSim (Wang et al., 2023) incorporates equivariance loss into vision-language models, significantly improving sensitivity to semantic changes in image-text pairs. Additionally, SoftCLIP (Gao et al., 2024) softens CLIP’s one-to-one constraint, enabling more flexible cross-modal alignment through fine-grained adjustments.
Compared with existing approaches, RankCLIP sets itself apart by fully leveraging the many-to-many relationships within each batch of text-image pairs, promoting learning from both matched and unmatched pairs with varying similarities by integrating in-modal and cross-modal list-wise ranking consistencies into the contrastive training objective. Crucially, RankCLIP diverges from existing models’ pair-wise training objective by adopting a global, list-wise optimization approach. In other words, it considers the rankings of all images and texts collectively within each batch, rather than focusing on pairwise similarities as seen in other methods.
3 RankCLIP
RankCLIP efficiently leverages the many-to-many relationships in real-world data by focusing on both matched and unmatched pairs. As shown in Fig. 2, it not only identifies if an image-text pair matches but also assesses their relative semantic similarities to other images and texts of both modalities in the dataset through self-supervised ranking consistency. Uniquely, RankCLIP employs a list-wise loss for training batches, distinguishing it from other methods that solely rely on pair-wise relationships, as discussed in §2.
3.1 Ranking Model Formulation
RankCLIP leverages the Plackett-Luce (PL) ranking model Plackett (1975); Luce (2005); Guiver and Snelson (2009) to estimate the probability distribution over rankings for every image-text pair , so that the consistency in their relative ordering with respect to a reference ranking can be measured. Specifically, for a given data pair, whether it is in-modal (image-image, text-text), or cross-modal (image-text), we calculate its in- or cross-modal cosine similarity to serve as the score when measuring the alignment of its ranking with respect to another reference ranking .
Following Plackett (1975), we first sort the reference ranking in a descending order to construct the optimal ranking , and assume that the ego ranking is sampled from . Thus the probability that item with score is ranked in the ego ranking from a set of items is the score of divided by the sum of scores for the items that have not been placed yet:
where denotes the set of items ranked before . In addition, we propose a decaying factor to scale the loss, so that the top-ranked items can obtain higher weights: Consequently, the probability of the entire ranking is the product of all individual placement probabilities:
RankCLIP’s objective is to maximize the consistency log-likelihood of the list ranking in one modality towards the reference ranking (from the same/in-modal and different/cross-modal data), which conveniently aligns with minimizing the negative log-likelihood loss:
where and can be in either modality.
3.2 Cross-modal Consistency Ranking
As illustrated by the green box in Fig. 2, RankCLIP utilizes secondary relationships between unmatched visual and textual representations by constructing a list-wise rank loss. This approach ensures that the semantic similarity rankings between one image and multiple texts align with those between one corresponding text and multiple images. For example, as shown in Fig. 1, from the dog perspective, the semantic distance between dog image and cat text is closer compared to the plane text. This relationship should also apply between the dog text and the cat, plane images. Mathematically, Eq. (3) can be specified as:
By optimizing Eq. (4), RankCLIP enhances its ability to bridge the semantic gap between modalities by leveraging nuanced unmatched correlations. This can also be viewed as learning a symmetric cosine-similarity matrix, further reinforcing semantic consistency across modalities.
3.3 In-modal Consistency Ranking
The pink box in Fig. 2 highlights the in-modal consistency component of the proposed rank loss. RankCLIP ensures semantic consistency within each modality – image to image and text to text – enhancing the use of secondary unmatched relationships as an optimization objective. The underlying principle is that similar images should correspond to similar texts. For example, in Fig. 1, from the dog image perspective, the cat image is the most similar, followed by the plane image. This relationship should hold true for their corresponding texts as well, where we utilize this to construct our and from Eq. (3). Mathematically, Eq. (3) can be specified as:
where and are the text and image batch embedding matrix, respectively. Via Eq. (6), the model can efficiently leverage the nuanced in-modal relationships to learn a richer and more structured semantic representation.
3.4 RankCLIP Loss
Combining both cross-modal and in-modal consistency with the traditional contrastive loss (more details in Appendix A), the complete rank loss is thus formulated as:
which is also depicted in Fig. 2. By supplementing the pairwise contrastive loss with cross-modal and in-modality ranking consistency loss, RankCLIP systematically organizes embeddings to fully exploit both global and fine-grained secondary unmatched relationships, which enhances the learning of more informative and accurate representations, better supporting downstream multi-modal tasks.
| CIFAR-10 | CIFAR-100 | ImageNet1K | |||||||
| Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | |
| CLIP | 36.35% | 70.28% | 85.02% | 12.22% | 24.93% | 33.56% | 12.08% | 21.86% | 27.48% |
| ALIP | 35.71% | 72.39% | 88.77% | 13.67% | 27.10% | 34.76% | 15.62% | 26.90% | 32.50% |
| RankCLIP | 37.03% (+1.87%) | 67.67% (-3.71%) | 83.09% (-2.27%) | 13.98% (+14.40%) | 27.70% (+11.11%) | 36.17% (+7.78%) | 17.02% (+40.89%) | 28.44% (+30.10%) | 33.99% (+23.69%) |
| ImageNetV2 | ImageNetSketch | ImageNet-A | ImageNet-R | |||||||||
| Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | |
| CLIP | 12.11% | 22.66% | 28.57% | 3.20% | 7.00% | 9.83% | 3.16% | 8.81% | 13.04% | 11.34% | 21.38% | 27.10% |
| ALIP | 15.62% | 27.34% | 32.82% | 5.10% | 10.37% | 14.01% | 3.53% | 9.14% | 13.61% | 14.25% | 25.74% | 32.43% |
| RankCLIP | 17.03% (+40.63%) | 28.60% (+26.21%) | 34.18% (+19.64%) | 5.82% (+81.88%) | 11.35% (+62.14%) | 14.87% (+51.27%) | 3.82% (+20.89%) | 9.16% (+3.97%) | 13.77% (+5.60%) | 15.74% (+38.80%) | 27.51% (+28.67%) | 34.36% (+26.79%) |
4 Experiments
4.1 Experimental Setup
Baselines.
The most direct baseline to RankCLIP is the original CLIP (Radford et al., 2021). In addition, to further demonstrate the superior performance of RankCLIP, we include ALIP (Yang et al., 2023), which leverages synthetic captions to enhance vision-language representation learning. More specifically, it employs a unique architecture that dynamically adjusts sample and pair weights to mitigate the impact of noisy or irrelevant data, which is quite orthogonal to our approach. The training procedures and parameters of all models are detailed in Appendix B.
Pretraining dataset.
Both baseline models, CLIP (Radford et al., 2021), ALIP (Yang et al., 2023) and the proposed RankCLIP are pretrained on the Conceptual Captions 3M (CC3M) dataset (Sharma et al., 2018), which contains around 3.3 million text-image pairs. Despite being much smaller than CLIP’s initial dataset (Ilharco et al., 2021), CC3M effectively supports the development of pretrained models with strong zero-shot capabilities and is widely used in existing language-image pretraining research (Carlini and Terzis, 2021; Li et al., 2021; Tejankar et al., 2021; Mu et al., 2022; Goel et al., 2022).
4.2 Zero-shot Classification
Zero-shot capability is one of the most significant improvements that CLIP achieves. Thus in this section, we first evaluate the zero-shot classification performance of CLIP, ALIP and the proposed RankCLIP. Following (Goel et al., 2022), we conduct our experiments on CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), and ImageNet1K (Deng et al., 2009; Russakovsky et al., 2015) dataset.
As shown in Table 1, RankCLIP consistently outperforms CLIP across CIFAR-10, CIFAR-100, and ImageNet1K datasets. Relative to CLIP, RankCLIP shows average improvements of +18.95%, +12.5%, and +9.73% in top-1, top-3 and top-5 metrics, respectively. Particularly on the more challenging ImageNet1K dataset, RankCLIP improves relative top-1 accuracy by +40.89% over CLIP, highlighting the effectiveness of the proposed ranking consistency in enhancing language-image alignment and understanding with the same amount of training data. The two cases where RankCLIP does not excel are the top-3 and top-5 accuracy on CIFAR-10. However, this is likely because CIFAR-10 with top-3 and top-5 metrics is much simpler, reducing the demand for a deeper model understanding.
Additionally, we observe that RankCLIP consistently outperforms ALIP, suggesting that our ranking consistency more effectively enhances text-image representations and alignments compared to the synthetic captions proposed in ALIP. Another trend we observe is that RankCLIP shows the most significant improvement in top-1 accuracy compared to top-3 and top-5. Considering the real-world emphasis on the topmost model output, RankCLIP is likely to offer considerable advantages in practical applications.
CIFAR-10 |
CIFAR-100 |
DTD |
FGVGAircraft |
Food101 |
GTSRB |
Imagenet1K |
OxfordPets |
SST2 |
STL10 |
SVHN |
Average | |
| CLIP | 72.40% | 48.43% | 49.89% | 26.10% | 48.59% | 65.20% | 77.49% | 49.74% | 53.71% | 83.59% | 44.80% | 56.37% |
| ALIP | 73.87% | 51.00% | 58.09% | 27.72% | 49.74% | 60.34% | 73.14% | 59.36% | 53.98% | 87.94% | 38.07% | 57.56% |
| RankCLIP | 72.54% (+0.20%) | 49.16% (+1.50%) | 53.24% (+6.71%) | 24.99% (-4.25%) | 47.11% (-3.05%) | 63.37% (-2.81%) | 86.40% (+11.50%) | 54.10% (+8.77%) | 54.09% (+3.00%) | 86.10% (+0.71%) | 43.30% (-3.35%) | 57.67% (+2.31%) |
| Flickr30K | MSCOCO | |||||||||||
| Text Retrieval | Image Retrieval | Text Retrieval | Image Retrieval | |||||||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R5 | R@10 | |
| CLIP | 84.00% | 88.70% | 91.0% | 8.70% | 16.90% | 21.20% | 82.06% | 85.24% | 87.82% | 5.04% | 12.98% | 18.32% |
| ALIP | 84.40% | 90.00% | 92.50% | 9.40% | 17.60% | 21.30% | 82.56% | 86.04% | 88.26% | 6.08% | 13.96% | 19.38% |
| RankCLIP | 84.10% (+0.12%) | 89.40% (+0.79%) | 91.90% (+0.99%) | 8.10% (-6.90%) | 16.40% (-2.96%) | 21.70% (+2.36%) | 82.90% (+1.02%) | 85.68% (+0.52%) | 88.00% (+0.20%) | 5.60% (+11.11%) | 13.20% (+1.69%) | 18.02% (-1.64%) |
4.3 Robustness to Distribution Shifts
To evaluate the robustness of RankCLIP under distribution shifts, we test it alongside CLIP and ALIP across four ImageNet variants, including ImageNetV2 (Recht et al., 2019), ImageNetSketch (Wang et al., 2019), ImageNet-A (Hendrycks et al., 2021b), and ImageNet-R (Hendrycks et al., 2021a), which are designed to assess resilience to different distribution shifts.
As shown in Table 2, RankCLIP outperforms both CLIP and ALIP consistently. Notably, relative to CLIP, RankCLIP’s accuracy improvements in shifted conditions are +45.55% (top-1), +30.24% (top-3), and +25.83% (top-5), surpassing its performance in standard settings (Table 1) of +40.89% (top-1), +30.1% (top-3), and +23.69% (top-5), indicating the even more superior performance in robustness under distribution shifts.
4.4 Linear Probing
We also evaluate whether the introduced ranking consistency retains its advantages when supplemented with additional in-domain supervision. Specifically, we use linear probing, where the pretrained encoders from CLIP, ALIP, and RankCLIP remain unchanged while a logistic regression classifier is trained on domain-specific datasets. We evaluate on a suite of 11 standard image classification datasets as our in-domain datasets, which include CIFAR-10, CIFAR-100, Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Fine-Grained Visual Classification of Aircraft (FGVG-Aircraft) (Maji et al., 2013), Food101 (Bossard et al., 2014), German Traffic Sign Detection Benchmark (GTSDB) (Stallkamp et al., 2012), ImageNet1K (Deng et al., 2009; Russakovsky et al., 2015), OxfordPets (Parkhi et al., 2012), Stanford Sentiment Treebank v2 (SST2) (Socher et al., 2013), STL-10 (Coates et al., 2011), and Street View House Numbers (SVHN) (Netzer et al., 2011) dataset.
Table 3 indicates that RankCLIP consistently outperforms CLIP, with relative improvements ranging from +0.2% to +11.5% and an average accuracy increase of +2.31%. When compared to ALIP, RankCLIP also shows better performance on average, though the gains are relatively modest.
4.5 Zero-shot Image-text Retrieval
In the final part of our experiments, we assess RankCLIP on zero-shot cross-modal retrieval tasks (image-to-text and text-to-image) using the Flickr30k (Plummer et al., 2015) and MSCOCO (Lin et al., 2014) datasets. As shown in Table 4, RankCLIP generally outperforms the two baseline methods, though improvements are less significant compared to earlier results in Table 1, 2 and 3. The relatively modest gains in retrieval tasks may stem from the complex requirements of discerning image-text similarities across varying resolutions and object details, a significant departure from the simpler demands of image classification tasks. Despite this, the overall improvement highlights RankCLIP’s advantage, thanks to the deeper insights provided by ranking consistency in the language-image training process.
5 Ablation Studies
| CIFAR-10 | CIFAR-100 | ImageNet1K | |||||||
| Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | |
| CLIP | 36.35% | 70.28% | 85.02% | 12.22% | 24.93% | 33.56% | 12.08% | 21.86% | 27.48% |
| RankCLIP | 37.03% | 67.67% | 83.09% | 13.98% | 27.70% | 36.17% | 17.02% | 28.44% | 33.99% |
| 37.47% | 69.89% | 84.53% | 13.89% | 27.34% | 35.90% | 16.66% | 27.63% | 33.15% | |
| 28.26% | 59.65% | 75.45% | 13.29% | 26.85% | 34.71% | 16.98% | 28.25% | 33.90% | |
5.1 Ablation on Loss Components
To further assess the effectiveness of the proposed ranking consistency, we developed two variants of RankCLIP: , focusing solely on cross-modal consistency with , and , emphasizing in-modal consistency with . Both models underwent the same pretraining as outlined in Appendix B and were tested in a zero-shot classification experiment on ImageNet1K as in §4.2. The results are shown in Table 5, with bold font indicating the best performance, and blue color representing the second best results. We can see that, while RankCLIP achieves the best performance, both and demonstrate notable improvements over CLIP. Interestingly, matches the performance of , highlighting the often-underestimated value of in-modal consistency in enhancing model effectiveness.
5.2 Ablation on Data Sizes
To evaluate the scalability of RankCLIP, we trained both CLIP and RankCLIP using 500K, 750K, 1M, and 3M text-image pairs from the CC3M dataset following the same procedures in §B. Fig. 3 left presents the zero-shot top-1 classification accuracy on ImageNet1K, where RankCLIP consistently outperforms CLIP. Notably, it shows a greater performance increase as dataset size grows from 1M to 3M pairs, suggesting RankCLIP’s superior scalability, a critical attribute for language-image pretraining. Furthermore, Fig. 3 right illustrates RankCLIP’s robustness across different dataset sizes. The horizontal axis shows the top-1 accuracy on standard ImageNet1K, and the vertical axis on ImageNet1K-R, with a black diagonal line () representing ideal robustness. Any deviation below this line indicates reduced robustness. RankCLIP consistently stays well above both the red baseline, which reflects typical in-distribution to out-of-distribution generalization (Miller et al., 2021), and close to the ideal line, demonstrating exceptional robustness to distribution shifts.
6 Analysis
6.1 Modality Gap
In this section, we analyze the modality gaps of CLIP and our proposed RankCLIP by visualizing 250 text-image pair embeddings, reduced to two dimensions using UMAP (McInnes et al., 2018), and complement this with a histogram of the gaps. Modality gap (Liang et al., 2022) refers to a geometric phenomenon observed in the representation spaces of multimodal models, where different data modalities (like images and texts) are embedded at a noticeable distance from each other, rather than being uniformly distributed as ideally expected. This gap, inherent from initialization and preserved during the contrastive learning process like in CLIP, poses a challenge in language-image pretraining by impacting joint data modeling and understanding. Recent studies (Srivastava and Sharma, 2024; Kumar and Marttinen, 2024; Oh et al., 2024) suggest that reducing this gap could enhance multimodal representations and downstream task performance. The results shown in Fig. 4 indicate that RankCLIP exhibits a significantly smaller modality gap than CLIP, demonstrating that our ranking consistency approach effectively enhances understanding of text-image semantics.
6.2 Alignment and Uniformity
Besides alleviating modality gap, it is also commonly believed that a successful contrastive learning method should as well ensure a broad and uniform distribution covering an hypersphere in space (Wang and Isola, 2020). These two goals, characterized as similarity and uniformity, can be assessed with alignment and uniformity scores, respectively. More specifically, following Goel et al. (2022) and notations defined in §3, we calculate the alignment score , and uniformity score to be:
where is the total number of text-image pairs. Essentially, represents the averaged cosine similarity between text and image embeddings, and averages the dissimilarity measures (exponentiated negative dot products) between all unique pairs of text-image embeddings in the dataset, quantifying how evenly these embeddings are distributed.
A high alignment score represents a strong correlation or similarity between pairs of text-image embeddings, indicating that the images and their textual descriptions are closely aligned in the embedding space. Conversely, a high uniformity score suggests that embeddings are not uniformly distributed; they may be clustering together or not utilizing the embedding space efficiently, which can indicate redundancy in the representations or a lack of diversity. A low uniformity score, on the other hand, suggests that the embeddings are well spread out across the space, indicating a diverse and efficient use of the embedding space, which is generally desirable for tasks like retrieval, where a wide coverage of possible queries are preferred.
As shown in Table 6, we observe that, although CLIP learns representations that are better aligned, as evidenced by its top-ranking alignment scores, these representations fail to achieve uniform distribution across the hypersphere, as highlighted by its significantly higher absolute uniformity scores. On the other hand, RankCLIP, along with two of its ablated version, and , presents much better balance between alignment and uniformity, which results in improved downstream task performance as illustrated in previous experiments as well as in the representative ZS-Top1 results in Table 6. We also find the results to be informative on a higher level where it indicates that optimizing contrastive learning towards single objective such as alignment or uniformity would not intuitively result in higher downstream task performance.
| CIFAR-10 | CIFAR-100 | ImageNet1K | |||||||
| ZS-Top1 | ZS-Top1 | ZS-Top1 | |||||||
| CLIP | 0.40 | -0.35 | 36.35% | 0.42 | -0.35 | 12.22% | 0.44 | -0.29 | 12.08% |
| RankCLIP | 0.23 | -0.17 | 37.03% | 0.26 | -0.16 | 13.98% | 0.33 | -0.11 | 17.02% |
| 0.24 | -0.16 | 37.47% | 0.26 | -0.15 | 13.89% | 0.32 | -0.10 | 16.66% | |
| 0.18 | -0.12 | 28.26% | 0.18 | -0.10 | 13.29% | 0.26 | -0.09 | 16.98% | |
7 Conclusion
In this paper, we introduce RankCLIP, a novel language-image pretraining method that integrates ranking consistency into the contrastive learning paradigm. RankCLIP aims to better understand the complex many-to-many relationships in diverse text-image pairs by optimizing a self-supervised, list-wise rank loss. Through extensive experiments, including zero-shot classification, robustness to distribution shifts, linear probing, and zero-shot image-text retrieval, RankCLIP not only enhances performance but also improves model robustness and semantic comprehension, outperforming the baseline CLIP and another state-of-the-art model ALIP. Our ablation studies and analyses further demonstrate and interpret the significance of each component of RankCLIP in boosting performance and understanding across modalities. We believe that the methodologies and principles of RankCLIP will inspire further research and lead to the development of models with a deeper understanding of the intricate interactions between visual and textual data.
Limitations
This work has two main limitations. First, due to computational constraints, we were not able to perform more baseline comparisons by pretraining additional existing methods. And second, we were unable to scale the pretraining process to include larger datasets, such as those used in OpenCLIP (Ilharco et al., 2021).
Appendix
Appendix A CLIP Preliminaries
CLIP (Radford et al., 2021) has been a prominent method for learning detailed multimodal representations through the alignment of images and texts. Given a set of image-text pairs, where denotes an image and is the corresponding text, the goal is to learn representations that map semantically similar images and texts closer in the embedding space, while dissimilar pairs are distanced apart. More specifically, the foundational CLIP model employs two encoders: an image encoder that processes raw images into visual embeddings and a text encoder which encodes textual data into text embeddings. Then both the text and visual features are projected to a latent space with identical dimension. Formally, the embeddings for a text-image pair are denoted as and , respectively. The embeddings are then normalized to lie on an unit hypersphere by enforcing -norm constraint:
so that the magnitude information is erased and only direction is preserved.
To align the image and text representations, a contrastive loss function, typically a variant of the InfoNCE loss (Oord et al., 2018), which optimizes the similarity of the matched pair against unmatched pairs, is utilized, i.e.:
where the first term contrasts images with the texts, the second term contrasts texts with the images, and denotes a temperature scaling parameter that adjusts the concentration of the distribution. The optimization of Eqn. (13) results in embeddings where the cosine similarity between matched image-text pairs is maximized in comparison to unmatched pairs, thus achieving the desired alignment in the joint embedding space.
Despite the efficacy of CLIP in learning correlated multimodal embeddings, it inherently relies on strict pairwise matched comparisons and fails to capture the more complex, fine-grained nature of semantic similarity within and across modalities that are generally treated as unmatched. This observation motivates the development of RankCLIP, which innovates beyond binary pairwise contrasts to consider holistic listwise consistency within and across modalities.
Appendix B Training Procedures
B.1 Implementation Details
For CLIP (Radford et al., 2021), we use the official implementation released by OpenAI111CLIP repository on GitHub: https://github.com/openai/CLIP.. And for ALIP (Yang et al., 2023), we also use the official implementation released by the paper authors222ALIP repository on GitHub: https://github.com/deepglint/ALIP.. As the proposed RankCLIP essentially shares the same model architecture (separate vision, text encoders, projection layer, and a classification head) as CLIP, we build upon the CLIP code repository for our model construction333RankCLIP repository will be released upon acceptance.. We set the scaling parameters for cross-modal () and in-modal () ranking consistency to 1/16 and 1/16 respectively throughout all the experiments unless otherwise noted. All CLIP, ALIP and RankCLIP models are initialized from scratch without loading any existing weights. And the embedding sizes for both modalities all project to 1024 across the three models.
B.2 Training Parameters
Following CLIP (Radford et al., 2021), we adopt the ResNet-50 (He et al., 2016) and transformer architectures (Devlin et al., 2018) for image and text encoding, respectively. Training is conducted from scratch over 64 epochs using a single NVIDIA A100 GPU, with a batch size of 512, an initial learning rate of 0.0005 employing cosine scheduling, and 10,000 warm-up steps.

### v3

@@
 Code and model checkpoints are available at https://github.com/Jam1ezhang/RankCLIP.
-RankCLIP: Ranking-Consistent Language-Image Pretraining
-Yiming Zhang††thanks: These authors contribute equally.
-Correspondence to Zenghui Ding <dingzenghui@iim.ac.cn>.
-Zhuokai Zhao Zhaorun Chen2 Zhili Feng3 Zenghui Ding1 Yining Sun1 1HFIPS, Chinese Academy of Sciences 2University of Chicago 3Carnegie Mellon University 4University of Science and Technology of China
 1 Introduction
-In the realm of computer vision (CV) (Voulodimos et al., 2018), natural language processing (NLP) (Chowdhary and Chowdhary, 2020), and multimodal deep learning (Jabeen et al., 2023; Zhao et al., 2023), the alignment between visual and textual modalities (Singh et al., 2022; Chen et al., 2024a) has emerged as a cornerstone for downstream applications, ranging from image captioning (Ghandi et al., 2023) to zero-shot classification (Pourpanah et al., 2022).
-Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021) marks a significant advancement in this field, demonstrating incredible performance from training on large amounts of text-image pairs to create self-supervised models that understand (Hendrycks et al., 2021a, b; Chen et al., 2024b) and generate (Ramesh et al., 2021; Crowson et al., 2022) descriptions of visual contents.
+In the realm of computer vision (CV) [55], natural language processing (NLP) [7], and multimodal deep learning [25, 60, 4], the alignment between visual and textual modalities [49, 5] has emerged as a cornerstone for downstream applications, ranging from image captioning [18] to zero-shot classification [42].
+Contrastive Language-Image Pretraining (CLIP) [43] marks a significant advancement in this field, demonstrating incredible performance from training on large amounts of text-image pairs to create self-supervised models that understand [22, 23, 6] and generate [44, 11] descriptions of visual contents.
 Following the success of this contrastive learning paradigm, many recent works have been developed upon the original CLIP.
-More specifically, these enhancements focus on optimizing data efficiency through intrinsic supervision (Li et al., 2021), as well as improving downstream performance via cross-modal late interaction (Yao et al., 2021), hierarchical feature alignment (Gao et al., 2022), geometric consistency regularization (Goel et al., 2022), additional learning (Mu et al., 2022), adaptive loss (Yang et al., 2023), hierarchy-aware attentions (Geng et al., 2023), and softer cross-modal alignment (Gao et al., 2024).
-Despite the improvements, these methods often have reliance on strict pairwise, cross-modal, and one-to-one mappings between images and texts, overlooking the actual many-to-many relationships that exist both cross-modal and in-modal in real-world data (Chun, 2023).
+More specifically, these enhancements focus on optimizing data efficiency through intrinsic supervision [28], as well as improving downstream performance via cross-modal late interaction [59], hierarchical feature alignment [15], geometric consistency regularization [19], additional learning [36], adaptive loss [58], hierarchy-aware attentions [17], and softer cross-modal alignment [16].
+Despite the improvements, these methods often have reliance on strict pairwise, cross-modal, and one-to-one mappings between images and texts, overlooking the actual many-to-many relationships that exist both cross-modal and in-modal in real-world data [8].
 For example, as shown in Fig.
 1, while pretrained models like CLIP can correctly classify dog, cat and airplane, they do not necessarily learn that dog and cat are more close to each other than dog and airplane, in terms of both in-modal (dog text is more similar to cat text than to airplane text) and cross-modal (dog text is more matched to cat image than to airplane image) similarities.
-Because it is rooted from the current contrastive loss that only the correct pairs are optimized while the rest of the unmatched pairs are treated the same, resulting in a large amount of information not used and unknown to the model during and after the training process.
+Since it is rooted from the current contrastive loss that only the correct pairs are optimized while the rest of the unmatched pairs are treated the same, a large amount of information not used and unknown to the model during and after the training process.
 Recognizing the complex many-to-many relationships as well as the rich information contained within both in-modal and cross-modal data, we introduce Ranking-Consistent Language-Image Pretraining, (RankCLIP), which employs ranking consistency to learn and optimize similarity levels both between (cross-modal) and within (in-modal) the text-image pairs.
@@
 Ranking consistency is conveniently modeled as an additional loss term to the traditional contrastive loss, requiring no extra external modules.
-It acts as a plug-and-play improvement for many existing methods, including those focusing on data-efficiency (Li et al., 2021), potentially boosting performance in both efficiency and effectiveness.
-The main contributions of this paper are: (1) RankCLIP, a novel contrastive pretraining method that uses ranking consistency to exploit the many-to-many relationships within data, thereby enhancing performance in downstream tasks such as zero-shot classification and retrieval accuracy; and (2) through comprehensive experiments conducted on multiple datasets, we demonstrate the superior effectiveness of RankCLIP in improving pretraining model performance without requiring any additional data or computational resources.
+It acts as a plug-and-play improvement for many existing methods, including those focusing on data-efficiency [28], potentially boosting performance in both efficiency and effectiveness.
+The main contributions of this paper are: 1) RankCLIP, a novel contrastive pretraining method that uses ranking consistency to exploit the many-to-many relationships within data, thereby enhancing performance in downstream tasks such as zero-shot classification and retrieval accuracy; and 2) through comprehensive experiments conducted on multiple datasets, we demonstrate the superior effectiveness of RankCLIP in improving pretraining model performance without requiring any additional data or computational resources.
 2 Related Work
-Vision-language pretraining has witnessed significant advancements over the past years (Chen et al., 2023; Du et al., 2022; Long et al., 2022).
-Models such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021) and FLAVA (Singh et al., 2022) have pioneered the contrastive learning paradigm applied with text-image pairs, showcasing remarkable performance and robustness in downstream tasks.
+Vision-language pretraining has witnessed significant advancements over the past years [3, 14, 31].
+Models such as CLIP [43], ALIGN [26] and FLAVA [49] have pioneered the contrastive learning paradigm applied with text-image pairs, showcasing remarkable performance and robustness in downstream tasks.
 Many follow-up works, mostly built upon CLIP, have been proposed since then.
 Li et al.
-(2021) introduced DeCLIP, improving zero-shot performance through intrinsic supervision.
-FILIP (Yao et al., 2021) advances CLIP’s alignment between image patches and text with a cross-modal interaction mechanism.
+[28] introduced DeCLIP, improving zero-shot performance through intrinsic supervision.
+FILIP [59] advances CLIP’s alignment between image patches and text with a cross-modal interaction mechanism.
 Gao et al.
-(2022) developed PyramidCLIP, using hierarchical feature alignment to boost model efficiency and performance.
-Additionally, SLIP (Mu et al., 2022) merges self-supervised learning with CLIP pre-training for improved visual representation and accuracy.
+[15] developed PyramidCLIP, using hierarchical feature alignment to boost model efficiency and performance.
+Additionally, SLIP [36] merges self-supervised learning with CLIP pre-training for improved visual representation and accuracy.
 Goel et al.
-(2022) introduced CyCLIP, augmenting CLIP with geometric consistency regularizers to enhance robustness and performance under varied conditions.
+[19] introduced CyCLIP, augmenting CLIP with geometric consistency regularizers to enhance robustness and performance under varied conditions.
 Recently, Yang et al.
-(2023) introduced ALIP, an adaptive pre-training model that enhances language-image alignment using raw text and synthetic captions with dynamic adjustments.
-HiCLIP (Geng et al., 2023) refines CLIP by adding hierarchy-aware attentions to uncover semantic hierarchies in images and texts.
-EqSim (Wang et al., 2023) incorporates equivariance loss into vision-language models, significantly improving sensitivity to semantic changes in image-text pairs.
-Additionally, SoftCLIP (Gao et al., 2024) softens CLIP’s one-to-one constraint, enabling more flexible cross-modal alignment through fine-grained adjustments.
+[58] introduced ALIP, an adaptive pre-training model that enhances language-image alignment using raw text and synthetic captions with dynamic adjustments.
+HiCLIP [17] refines CLIP by adding hierarchy-aware attentions to uncover semantic hierarchies in images and texts.
+EqSim [57] incorporates equivariance loss into vision-language models, significantly improving sensitivity to semantic changes in image-text pairs.
+Additionally, SoftCLIP [16] softens CLIP’s one-to-one constraint, enabling more flexible cross-modal alignment through fine-grained adjustments.
 Compared with existing approaches, RankCLIP sets itself apart by fully leveraging the many-to-many relationships within each batch of text-image pairs, promoting learning from both matched and unmatched pairs with varying similarities by integrating in-modal and cross-modal list-wise ranking consistencies into the contrastive training objective.
@@
 RankCLIP efficiently leverages the many-to-many relationships in real-world data by focusing on both matched and unmatched pairs.
-As shown in Fig.
+As in Fig.
 2, it not only identifies if an image-text pair matches but also assesses their relative semantic similarities to other images and texts of both modalities in the dataset through self-supervised ranking consistency.
@@
 3.1 Ranking Model Formulation
-RankCLIP leverages the Plackett-Luce (PL) ranking model Plackett (1975); Luce (2005); Guiver and Snelson (2009) to estimate the probability distribution over rankings for every image-text pair , so that the consistency in their relative ordering with respect to a reference ranking can be measured.
+RankCLIP leverages the Plackett-Luce (PL) ranking model [41, 32, 20] to estimate the probability distribution over rankings for every image-text pair , so that the consistency in their relative ordering with respect to a reference ranking can be measured.
 Specifically, for a given data pair, whether it is in-modal (image-image, text-text), or cross-modal (image-text), we calculate its in- or cross-modal cosine similarity to serve as the score when measuring the alignment of its ranking with respect to another reference ranking .
-Following Plackett (1975), we first sort the reference ranking in a descending order to construct the optimal ranking , and assume that the ego ranking is sampled from .
-Thus the probability that item with score is ranked in the ego ranking from a set of items is the score of divided by the sum of scores for the items that have not been placed yet:
-where denotes the set of items ranked before .
-In addition, we propose a decaying factor to scale the loss, so that the top-ranked items can obtain higher weights: Consequently, the probability of the entire ranking is the product of all individual placement probabilities:
+Following [41], we first sort the reference ranking in a descending order to construct the optimal ranking , and assume that the ego ranking is sampled from .
+The probability that item with score is ranked in the ego ranking from a set of items is the score of divided by the sum of scores for the items that have not been placed yet:
+where denotes the set of items ranked before . Consequently, the probability of the entire ranking is the product of all individual placement probabilities:
 RankCLIP’s objective is to maximize the consistency log-likelihood of the list ranking in one modality towards the reference ranking (from the same/in-modal and different/cross-modal data), which conveniently aligns with minimizing the negative log-likelihood loss:
-where and can be in either modality.
 3.2 Cross-modal Consistency Ranking
@@
 This can also be viewed as learning a symmetric cosine-similarity matrix, further reinforcing semantic consistency across modalities.
+ImageNet1K | MSCOCO | ||||||||
+Image-to-Text Retrieval | Text-to-Image Retrieval | ||||||||
+Top-1 | Top-3 | Top-5 | Recall@1 | Recall@5 | Recall@10 | Recall@1 | Recall@5 | Recall@10 | |
+CLIP [43] | 9.06% | 16.94% | 21.63% | 6.68% | 18.36% | 26.94% | 3.70% | 9.74% | 14.04% |
+CyCLIP [19] | 9.40% | 17.32% | 21.72% | 6.50% | 19.34% | 29.14% | 3.72% | 11.16% | 16.06% |
+ALIP [58] | 9.71% | 18.31% | 23.07% | 6.04% | 18.04% | 26.92% | 3.70% | 10.22% | 14.38% |
+RankCLIP | 10.16% | 19.57% | 24.01% | 7.18% | 19.46% | 28.48% | 3.74% | 10.28% | 14.18% |
 3.3 In-modal Consistency Ranking
@@
 3.4 RankCLIP Loss
-Combining both cross-modal and in-modal consistency with the traditional contrastive loss (more details in Appendix A), the complete rank loss is thus formulated as:
+Combining both cross-modal and in-modal consistency with the traditional contrastive loss (more details in Appendix B), the complete rank loss is thus formulated as:
 which is also depicted in Fig.
 2.
-By supplementing the pairwise contrastive loss with cross-modal and in-modality ranking consistency loss, RankCLIP systematically organizes embeddings to fully exploit both global and fine-grained secondary unmatched relationships, which enhances the learning of more informative and accurate representations, better supporting downstream multi-modal tasks.
-| CIFAR-10 | CIFAR-100 | ImageNet1K | |||||||
-| Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | |
-| CLIP | 36.35% | 70.28% | 85.02% | 12.22% | 24.93% | 33.56% | 12.08% | 21.86% | 27.48% |
-| ALIP | 35.71% | 72.39% | 88.77% | 13.67% | 27.10% | 34.76% | 15.62% | 26.90% | 32.50% |
-| RankCLIP | 37.03% (+1.87%) | 67.67% (-3.71%) | 83.09% (-2.27%) | 13.98% (+14.40%) | 27.70% (+11.11%) | 36.17% (+7.78%) | 17.02% (+40.89%) | 28.44% (+30.10%) | 33.99% (+23.69%) |
-| ImageNetV2 | ImageNetSketch | ImageNet-A | ImageNet-R | |||||||||
-| Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | |
-| CLIP | 12.11% | 22.66% | 28.57% | 3.20% | 7.00% | 9.83% | 3.16% | 8.81% | 13.04% | 11.34% | 21.38% | 27.10% |
-| ALIP | 15.62% | 27.34% | 32.82% | 5.10% | 10.37% | 14.01% | 3.53% | 9.14% | 13.61% | 14.25% | 25.74% | 32.43% |
-| RankCLIP | 17.03% (+40.63%) | 28.60% (+26.21%) | 34.18% (+19.64%) | 5.82% (+81.88%) | 11.35% (+62.14%) | 14.87% (+51.27%) | 3.82% (+20.89%) | 9.16% (+3.97%) | 13.77% (+5.60%) | 15.74% (+38.80%) | 27.51% (+28.67%) | 34.36% (+26.79%) |
+By supplementing the pairwise contrastive loss with cross-modal and in-modality ranking consistency loss, RankCLIP systematically organizes embeddings to fully exploit both global and fine-grained unmatched relationships, which enhances the learning of more informative and accurate representations, better supporting downstream multi-modal tasks.
+3.5 Training Recipe on Selecting and
+In the early stage of pre-training, rank consistency is highly unstable due to random initialization.
+Overemphasizing ranking consistency at this stage can impede the optimization of the embedding space.
+To address this, we gradually increase the weights and of the ranking loss as training progresses.
+Specifically, we have:
+where and denote the current training epoch and total number of epoch, respectively. The full RankCLIP framework is outlined in Algorithm 1.
 4 Experiments
@@
 Baselines.
-The most direct baseline to RankCLIP is the original CLIP (Radford et al., 2021).
-In addition, to further demonstrate the superior performance of RankCLIP, we include ALIP (Yang et al., 2023), which leverages synthetic captions to enhance vision-language representation learning.
-More specifically, it employs a unique architecture that dynamically adjusts sample and pair weights to mitigate the impact of noisy or irrelevant data, which is quite orthogonal to our approach.
-The training procedures and parameters of all models are detailed in Appendix B.
-Pretraining dataset.
-Both baseline models, CLIP (Radford et al., 2021), ALIP (Yang et al., 2023) and the proposed RankCLIP are pretrained on the Conceptual Captions 3M (CC3M) dataset (Sharma et al., 2018), which contains around 3.3 million text-image pairs.
-Despite being much smaller than CLIP’s initial dataset (Ilharco et al., 2021), CC3M effectively supports the development of pretrained models with strong zero-shot capabilities and is widely used in existing language-image pretraining research (Carlini and Terzis, 2021; Li et al., 2021; Tejankar et al., 2021; Mu et al., 2022; Goel et al., 2022).
-4.2 Zero-shot Classification
-Zero-shot capability is one of the most significant improvements that CLIP achieves.
-Thus in this section, we first evaluate the zero-shot classification performance of CLIP, ALIP and the proposed RankCLIP.
-Following (Goel et al., 2022), we conduct our experiments on CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), and ImageNet1K (Deng et al., 2009; Russakovsky et al., 2015) dataset.
-As shown in Table 1, RankCLIP consistently outperforms CLIP across CIFAR-10, CIFAR-100, and ImageNet1K datasets.
-Relative to CLIP, RankCLIP shows average improvements of +18.95%, +12.5%, and +9.73% in top-1, top-3 and top-5 metrics, respectively.
-Particularly on the more challenging ImageNet1K dataset, RankCLIP improves relative top-1 accuracy by +40.89% over CLIP, highlighting the effectiveness of the proposed ranking consistency in enhancing language-image alignment and understanding with the same amount of training data.
-The two cases where RankCLIP does not excel are the top-3 and top-5 accuracy on CIFAR-10.
-However, this is likely because CIFAR-10 with top-3 and top-5 metrics is much simpler, reducing the demand for a deeper model understanding.
-Additionally, we observe that RankCLIP consistently outperforms ALIP, suggesting that our ranking consistency more effectively enhances text-image representations and alignments compared to the synthetic captions proposed in ALIP.
-Another trend we observe is that RankCLIP shows the most significant improvement in top-1 accuracy compared to top-3 and top-5.
-Considering the real-world emphasis on the topmost model output, RankCLIP is likely to offer considerable advantages in practical applications.
+The most direct baseline to RankCLIP is the original CLIP [43].
+To further demonstrate the superior performance of RankCLIP, we include CyCLIP [19], which introduces cyclic consistency constraints to enforce more robust alignment between visual and textual representations, improving generalization and semantic coherence.
+We also include ALIP [58], which leverages synthetic captions to enhance vision-language representation learning.
+More specifically, it employs a unique architecture that dynamically adjusts sample and pair weights to mitigate the impact of noisy or irrelevant data, making its approach complementary to ours.
+The training procedures and parameters for all models are detailed in Appendix A.
+Data.
+All approaches are pretrained on the Conceptual Captions 3M (CC3M) dataset [48], which contains approximately 3.3 million text-image pairs.
+Although significantly smaller than CLIP’s original dataset [24], CC3M remains a standard benchmark for vision-language pretraining, enabling strong zero-shot performance [2, 28, 53, 36, 19].
+As discussed in §5.2, we also train CLIP and RankCLIP on a larger 15M-image subset of YFCC100M [54] (YFCC15M) to study the effect of scaling up the dataset size.
+ImageNetV2-Matched | ImageNetV2-Threshold | ImageNetV2-Top | ImageNet-R | |||||||||
+Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | |
+CLIP [43] | 7.53% | 14.99% | 19.61% | 8.89% | 17.22% | 21.86% | 10.76% | 19.80% | 24.87% | 9.36% | 10.56% | 19.76% |
+CyCLIP [19] | 7.68% | 15.07% | 19.11% | 9.10% | 17.42% | 21.94% | 11.20% | 20.18% | 25.34% | 9.23% | 16.72% | 21.64% |
+ALIP [58] | 7.82% | 15.56% | 19.81% | 9.65% | 18.31% | 22.85% | 11.43% | 20.88% | 26.10% | 10.92% | 20.27% | 26.24% |
+RankCLIP | 9.01% | 16.95% | 21.12% | 10.32% | 19.31% | 24.13% | 12.31% | 22.11% | 27.17% | 11.34% | 20.88% | 26.94% |
 CIFAR-10 |
@@
 GTSRB |
-Imagenet1K |
 OxfordPets |
@@
 SVHN |
-Average | |
-| CLIP | 72.40% | 48.43% | 49.89% | 26.10% | 48.59% | 65.20% | 77.49% | 49.74% | 53.71% | 83.59% | 44.80% | 56.37% |
-| ALIP | 73.87% | 51.00% | 58.09% | 27.72% | 49.74% | 60.34% | 73.14% | 59.36% | 53.98% | 87.94% | 38.07% | 57.56% |
-| RankCLIP | 72.54% (+0.20%) | 49.16% (+1.50%) | 53.24% (+6.71%) | 24.99% (-4.25%) | 47.11% (-3.05%) | 63.37% (-2.81%) | 86.40% (+11.50%) | 54.10% (+8.77%) | 54.09% (+3.00%) | 86.10% (+0.71%) | 43.30% (-3.35%) | 57.67% (+2.31%) |
-| Flickr30K | MSCOCO | |||||||||||
-| Text Retrieval | Image Retrieval | Text Retrieval | Image Retrieval | |||||||||
-| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R5 | R@10 | |
-| CLIP | 84.00% | 88.70% | 91.0% | 8.70% | 16.90% | 21.20% | 82.06% | 85.24% | 87.82% | 5.04% | 12.98% | 18.32% |
-| ALIP | 84.40% | 90.00% | 92.50% | 9.40% | 17.60% | 21.30% | 82.56% | 86.04% | 88.26% | 6.08% | 13.96% | 19.38% |
-| RankCLIP | 84.10% (+0.12%) | 89.40% (+0.79%) | 91.90% (+0.99%) | 8.10% (-6.90%) | 16.40% (-2.96%) | 21.70% (+2.36%) | 82.90% (+1.02%) | 85.68% (+0.52%) | 88.00% (+0.20%) | 5.60% (+11.11%) | 13.20% (+1.69%) | 18.02% (-1.64%) |
+Average |
+CLIP [43] | 77.6% | 56.2% | 43.2% | 22.6% | 39.7% | 60.0% | 40.4% | 51.0% | 79.0% | 50.5% | 52.0% |
+CyCLIP [19] | 76.8% | 54.3% | 45.8% | 19.2% | 37.5% | 58.6% | 44.2% | 51.5% | 82.3% | 41.3% | 51.2% |
+ALIP [58] | 71.1% | 49.1% | 47.1% | 17.4% | 36.1% | 51.5% | 41.9% | 53.3% | 81.0% | 38.3% | 48.7% |
+RankCLIP | 78.4% | 56.6% | 42.4% | 23.4% | 40.2% | 60.6% | 40.6% | 53.4% | 79.6% | 47.7% | 52.3% |
+4.2 Zero-shot Classification
+We evaluate the zero-shot classification performance of CLIP [43], CyCLIP [19], ALIP [58], and RankCLIP on ImageNet1K [12, 46].
+As shown in Table 1, RankCLIP consistently outperforms CLIP, which highlight the effectiveness of ranking consistency in enhancing language-image alignment with the same training data.
+Compared to CyCLIP [19], which enforces cyclic consistency to improve semantic coherence, RankCLIP achieves higher accuracy across all metrics, suggesting that ranking consistency provides a more direct and effective regularization for representation learning.
+Additionally, RankCLIP surpasses ALIP [58], indicating that ranking consistency is a stronger alternative to synthetic caption-based supervision.
+Notably, RankCLIP shows the most significant improvement in top-1 accuracy, reinforcing its practical advantages where the highest-ranked prediction is most critical.
+4.3 Zero-shot Cross-modal Retrieval
+We further evaluate RankCLIP on zero-shot cross-modal retrieval tasks, including image-to-text and text-to-image retrieval, using the MSCOCO [30] dataset.
+As shown in Table 1, RankCLIP outperforms all baselines, though the improvements are less pronounced compared to zero-shot classification.
+This smaller margin of improvement may be attributed to the increased complexity of retrieval tasks, which require fine-grained image-text alignment across varying resolutions and object details – challenges distinct from the more direct pattern recognition in classification.
 4.4 Robustness to Distribution Shifts
-To evaluate the robustness of RankCLIP under distribution shifts, we test it alongside CLIP and ALIP across four ImageNet variants, including ImageNetV2 (Recht et al., 2019), ImageNetSketch (Wang et al., 2019), ImageNet-A (Hendrycks et al., 2021b), and ImageNet-R (Hendrycks et al., 2021a), which are designed to assess resilience to different distribution shifts.
-As shown in Table 2, RankCLIP outperforms both CLIP and ALIP consistently.
-Notably, relative to CLIP, RankCLIP’s accuracy improvements in shifted conditions are +45.55% (top-1), +30.24% (top-3), and +25.83% (top-5), surpassing its performance in standard settings (Table 1) of +40.89% (top-1), +30.1% (top-3), and +23.69% (top-5), indicating the even more superior performance in robustness under distribution shifts.
+To evaluate the robustness of RankCLIP under distribution shifts, we test all approaches on three variants of ImageNetV2 [45] and ImageNet-R [22], which assess resilience to different real-world deviations from ImageNet1K.
+As shown in Table 2, RankCLIP consistently outperforms all baselines, achieving the highest accuracy across all datasets.
+These results indicate that RankCLIP not only improves standard zero-shot classification but also enhances adaptation to real-world distribution shifts.
 4.5 Linear Probing
-We also evaluate whether the introduced ranking consistency retains its advantages when supplemented with additional in-domain supervision.
-Specifically, we use linear probing, where the pretrained encoders from CLIP, ALIP, and RankCLIP remain unchanged while a logistic regression classifier is trained on domain-specific datasets.
-We evaluate on a suite of 11 standard image classification datasets as our in-domain datasets, which include CIFAR-10, CIFAR-100, Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Fine-Grained Visual Classification of Aircraft (FGVG-Aircraft) (Maji et al., 2013), Food101 (Bossard et al., 2014), German Traffic Sign Detection Benchmark (GTSDB) (Stallkamp et al., 2012), ImageNet1K (Deng et al., 2009; Russakovsky et al., 2015), OxfordPets (Parkhi et al., 2012), Stanford Sentiment Treebank v2 (SST2) (Socher et al., 2013), STL-10 (Coates et al., 2011), and Street View House Numbers (SVHN) (Netzer et al., 2011) dataset.
-Table 3 indicates that RankCLIP consistently outperforms CLIP, with relative improvements ranging from +0.2% to +11.5% and an average accuracy increase of +2.31%.
-When compared to ALIP, RankCLIP also shows better performance on average, though the gains are relatively modest.
-4.5 Zero-shot Image-text Retrieval
-In the final part of our experiments, we assess RankCLIP on zero-shot cross-modal retrieval tasks (image-to-text and text-to-image) using the Flickr30k (Plummer et al., 2015) and MSCOCO (Lin et al., 2014) datasets.
-As shown in Table 4, RankCLIP generally outperforms the two baseline methods, though improvements are less significant compared to earlier results in Table 1, 2 and 3.
-The relatively modest gains in retrieval tasks may stem from the complex requirements of discerning image-text similarities across varying resolutions and object details, a significant departure from the simpler demands of image classification tasks.
-Despite this, the overall improvement highlights RankCLIP’s advantage, thanks to the deeper insights provided by ranking consistency in the language-image training process.
+We further assess whether the advantages of ranking consistency persist when supplemented with in-domain supervision.
+Specifically, we apply linear probing, where pretrained encoders remain fixed while a logistic regression classifier is trained on domain-specific datasets.
+We evaluate on 10 standard image classification benchmarks, including CIFAR-10, CIFAR-100, DTD [9], FGVG-Aircraft [33], Food101 [1], GTSDB [52], OxfordPets [40], SST2 [50], STL-10 [10], and SVHN [37].
+As shown in Table 3, RankCLIP achieves the highest average accuracy, demonstrating that ranking consistency enhances generalization even with additional in-domain supervision.
 5 Ablation Studies
-| CIFAR-10 | CIFAR-100 | ImageNet1K | |||||||
-| Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | |
-| CLIP | 36.35% | 70.28% | 85.02% | 12.22% | 24.93% | 33.56% | 12.08% | 21.86% | 27.48% |
-| RankCLIP | 37.03% | 67.67% | 83.09% | 13.98% | 27.70% | 36.17% | 17.02% | 28.44% | 33.99% |
-| 37.47% | 69.89% | 84.53% | 13.89% | 27.34% | 35.90% | 16.66% | 27.63% | 33.15% | |
-| 28.26% | 59.65% | 75.45% | 13.29% | 26.85% | 34.71% | 16.98% | 28.25% | 33.90% | |
-5.1 Ablation on Loss Components
-To further assess the effectiveness of the proposed ranking consistency, we developed two variants of RankCLIP: , focusing solely on cross-modal consistency with , and , emphasizing in-modal consistency with .
-Both models underwent the same pretraining as outlined in Appendix B and were tested in a zero-shot classification experiment on ImageNet1K as in §4.2.
-The results are shown in Table 5, with bold font indicating the best performance, and blue color representing the second best results.
-We can see that, while RankCLIP achieves the best performance, both and demonstrate notable improvements over CLIP.
-Interestingly, matches the performance of , highlighting the often-underestimated value of in-modal consistency in enhancing model effectiveness.
-5.2 Ablation on Data Sizes
-To evaluate the scalability of RankCLIP, we trained both CLIP and RankCLIP using 500K, 750K, 1M, and 3M text-image pairs from the CC3M dataset following the same procedures in §B.
+5.1 Different Weights of RankCLIP Loss
+In Eq.
+(8), we define the RankCLIP loss as a linear combination of the original contrastive loss and the in-modal and cross-modal ranking consistency losses, weighted by and , respectively.
+In §3.5, we introduce a training strategy that adaptively adjusts and at different stages of training.
+In this section, we analyze the impact of these weights and demonstrates that adaptive weighting further enhances RankCLIP ’s performance.
+Notice that all model variants follow the same pretraining setup detailed in Appendix A.
+As shown in Fig.
+3, RankCLIP outperforms CLIP even with fixed and , highlighting the effectiveness of ranking consistency.
+And the adaptive weighting strategy further boosts accuracy by preventing the under-utilization of ranking consistency at low weights and avoiding disruptions to contrastive learning at high weights.
+Method | Vision Backbone | ImageNet1K | MSCOCO | Linear Probing Avg. Acc. | |||||||
+Image-to-Text Retrieval | Text-to-Image Retrieval | ||||||||||
+Top-1 | Top-3 | Top-5 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | |||
+CLIP [43] | RN50 | 21.6% | 36.9% | 44.9% | 15.6% | 36.4% | 48.4% | 6.7% | 15.2% | 20.1% | 64.2% |
+RankCLIP | 30.9% | 49.4% | 57.6% | 19.5% | 42.6% | 54.8% | 7.5% | 16.2% | 21.6% | 68.9% | |
+CLIP [43] | ViT-B/32 | 20.7% | 35.0% | 42.4% | 11.9% | 29.4% | 40.8% | 5.1% | 12.9% | 17.9% | 60.7% |
+RankCLIP | 26.2% | 41.4% | 48.9% | 13.8% | 33.8% | 45.9% | 6.0% | 13.6% | 18.6% | 61.3% |
+5.2 Different Data Sizes
+To assess the scalability of RankCLIP, we trained both CLIP and RankCLIP on 500k, 750k, 1M, and 3M text-image pairs from CC3M, as well as 15M pairs from YFCC15M, following the procedure in Appendix A.
 Fig.
-3 left presents the zero-shot top-1 classification accuracy on ImageNet1K, where RankCLIP consistently outperforms CLIP.
-Notably, it shows a greater performance increase as dataset size grows from 1M to 3M pairs, suggesting RankCLIP’s superior scalability, a critical attribute for language-image pretraining.
-Furthermore, Fig.
-3 right illustrates RankCLIP’s robustness across different dataset sizes.
-The horizontal axis shows the top-1 accuracy on standard ImageNet1K, and the vertical axis on ImageNet1K-R, with a black diagonal line () representing ideal robustness.
-Any deviation below this line indicates reduced robustness.
-RankCLIP consistently stays well above both the red baseline, which reflects typical in-distribution to out-of-distribution generalization (Miller et al., 2021), and close to the ideal line, demonstrating exceptional robustness to distribution shifts.
+4 compares their performance on zero-shot top-1 classification accuracy for ImageNet1K (left) and averaged linear probing results (middle), where RankCLIP consistently outperforms CLIP.
+Full, non-averaged linear probing results are provided in Appendix 7.
+Notably, the performance gains of RankCLIP become more pronounced as dataset size scales from 1m to 15m, highlighting its superior scalability, which is critical for language-image pretraining.
+Fig.
+4 (right) further illustrates RankCLIP’s robustness across different dataset sizes.
+The horizontal axis represents top-1 accuracy on standard ImageNet1K, while the vertical axis shows accuracy on ImageNet1K-R.
+The black diagonal () denotes ideal robustness, with deviations below it indicating degradation under distribution shifts.
+RankCLIP remains well above the red baseline, which represents typical in-distribution to out-of-distribution generalization [35], and stays close to the ideal line, demonstrating strong robustness.
+5.3 Different Backbones: RN50 vs. ViT
+We compare RankCLIP and CLIP across ResNet-50 (RN50) and ViT-B/32 backbones to assess its generalization.
+As shown in Table 4, RankCLIP consistently outperforms CLIP across all tasks.
+Specifically, for zero-shot classification on ImageNet1K, RankCLIP improves top-1 accuracy by +9.3% with RN50 and +5.5% with ViT-B/32, demonstrating stronger feature learning.
+In cross-modal retrieval, RankCLIP achieves +3.9% in image-to-text and +0.8% in text-to-image with RN50, with smaller but consistent gains for ViT-B/32.
+RankCLIP also improves linear probing accuracy by +4.7% with RN50 and +0.6% with ViT-B/32, confirming its advantage in representation learning.
+While both architectures benefit, RN50 sees the largest gains, suggesting that ranking consistency particularly enhances hierarchical feature extraction in CNNs.
 6 Analysis
 6.1 Modality Gap
-In this section, we analyze the modality gaps of CLIP and our proposed RankCLIP by visualizing 250 text-image pair embeddings, reduced to two dimensions using UMAP (McInnes et al., 2018), and complement this with a histogram of the gaps.
-Modality gap (Liang et al., 2022) refers to a geometric phenomenon observed in the representation spaces of multimodal models, where different data modalities (like images and texts) are embedded at a noticeable distance from each other, rather than being uniformly distributed as ideally expected.
-This gap, inherent from initialization and preserved during the contrastive learning process like in CLIP, poses a challenge in language-image pretraining by impacting joint data modeling and understanding.
-Recent studies (Srivastava and Sharma, 2024; Kumar and Marttinen, 2024; Oh et al., 2024) suggest that reducing this gap could enhance multimodal representations and downstream task performance.
-The results shown in Fig.
-4 indicate that RankCLIP exhibits a significantly smaller modality gap than CLIP, demonstrating that our ranking consistency approach effectively enhances understanding of text-image semantics.
+We analyze the modality gaps of CLIP and RankCLIP by visualizing 250 text-image pair embeddings, reduced to two dimensions using UMAP [34], and presenting a histogram of the gaps.
+The modality gap [29] refers to the separation between text and image embeddings in multimodal models, hindering joint representation learning.
+This gap, inherent from initialization and reinforced by contrastive learning in CLIP, challenges effective language-image modeling.
+Recent studies [51, 27, 38] suggest that reducing this gap improves multimodal representations and downstream performance.
+As shown in Fig.
+6, RankCLIP exhibits a significantly smaller modality gap than CLIP, demonstrating that our ranking consistency approach effectively enhances text-image alignment.
 6.2 Alignment and Uniformity
-Besides alleviating modality gap, it is also commonly believed that a successful contrastive learning method should as well ensure a broad and uniform distribution covering an hypersphere in space (Wang and Isola, 2020).
-These two goals, characterized as similarity and uniformity, can be assessed with alignment and uniformity scores, respectively.
-More specifically, following Goel et al.
-(2022) and notations defined in §3, we calculate the alignment score , and uniformity score to be:
-where is the total number of text-image pairs.
-Essentially, represents the averaged cosine similarity between text and image embeddings, and averages the dissimilarity measures (exponentiated negative dot products) between all unique pairs of text-image embeddings in the dataset, quantifying how evenly these embeddings are distributed.
-A high alignment score represents a strong correlation or similarity between pairs of text-image embeddings, indicating that the images and their textual descriptions are closely aligned in the embedding space.
-Conversely, a high uniformity score suggests that embeddings are not uniformly distributed; they may be clustering together or not utilizing the embedding space efficiently, which can indicate redundancy in the representations or a lack of diversity.
-A low uniformity score, on the other hand, suggests that the embeddings are well spread out across the space, indicating a diverse and efficient use of the embedding space, which is generally desirable for tasks like retrieval, where a wide coverage of possible queries are preferred.
-As shown in Table 6, we observe that, although CLIP learns representations that are better aligned, as evidenced by its top-ranking alignment scores, these representations fail to achieve uniform distribution across the hypersphere, as highlighted by its significantly higher absolute uniformity scores.
-On the other hand, RankCLIP, along with two of its ablated version, and , presents much better balance between alignment and uniformity, which results in improved downstream task performance as illustrated in previous experiments as well as in the representative ZS-Top1 results in Table 6.
-We also find the results to be informative on a higher level where it indicates that optimizing contrastive learning towards single objective such as alignment or uniformity would not intuitively result in higher downstream task performance.
-| CIFAR-10 | CIFAR-100 | ImageNet1K | |||||||
-| ZS-Top1 | ZS-Top1 | ZS-Top1 | |||||||
-| CLIP | 0.40 | -0.35 | 36.35% | 0.42 | -0.35 | 12.22% | 0.44 | -0.29 | 12.08% |
-| RankCLIP | 0.23 | -0.17 | 37.03% | 0.26 | -0.16 | 13.98% | 0.33 | -0.11 | 17.02% |
-| 0.24 | -0.16 | 37.47% | 0.26 | -0.15 | 13.89% | 0.32 | -0.10 | 16.66% | |
-| 0.18 | -0.12 | 28.26% | 0.18 | -0.10 | 13.29% | 0.26 | -0.09 | 16.98% | |
+Beyond reducing modality gap, effective contrastive learning should ensure a broad and uniform distribution over a hypersphere [56].
+These objectives – similarity and uniformity – are quantified by alignment and uniformity scores, respectively.
+Following Goel et al.
+[19] and the notations in §3, we compute the alignment score and uniformity score as:
+where is the number of text-image pairs.
+captures the average cosine similarity between corresponding text and image embeddings, while measures how evenly embeddings are spread across the space.
+High alignment indicates strong correlation between paired embeddings, whereas low uniformity suggests diverse and efficient embedding distribution—desirable for tasks like retrieval.
+As shown in Table 5, CLIP achieves stronger alignment but suffers from poor uniformity, leading to redundant representations.
+On the other hand, RankCLIP along with two of its ablated version, and presents much better balance between alignment and uniformity.
+These results further suggest that optimizing solely for alignment or uniformity does not necessarily translate to better task performance.
+CIFAR-10 | CIFAR-100 | ImageNet1K | ||||
+CLIP | 0.28 | -0.19 | 0.28 | -0.18 | 0.33 | -0.19 |
+RankCLIP | 0.23 | -0.14 | 0.26 | -0.13 | 0.29 | -0.13 |
+0.23 | -0.13 | 0.24 | -0.12 | 0.29 | -0.14 | |
+0.25 | -0.15 | 0.28 | -0.14 | 0.32 | -0.17 |
+6.3 Qualitative Examples
+Class activation maps.
+To further examine the effects of ranking consistency, we visualize class activation maps (CAMs) [47] for RankCLIP and CLIP in Fig.
+7.
+The results show that RankCLIP consistently attends to more semantically relevant regions in the images.
+For example, when given the caption ‘An airport filled with planes sitting on tarmacs’, CLIP mistakenly highlights surrounding areas, whereas RankCLIP focuses precisely on the planes.
+Similar improvements are observed across other examples, demonstrating that RankCLIP better aligns textual descriptions with visual concepts.
+This suggests that ranking consistency enhances fine-grained feature learning, leading to more localized and accurate visual grounding.
+Text-to-image retrieval.
+Fig.
+5 compares text-to-image retrieval results from RankCLIP and CLIP.
+Given a query, we display the top-ranked images retrieved by each model.
+RankCLIP consistently retrieves more semantically aligned images, even beyond the correct reference image.
+For instance, in the example of ’A cute cat laying down in a sink’, correctly identifies a cat in a sink, whereas CLIP misidentifies it due to the visual similarity between sinks and toilets.
+This demonstrates RankCLIP ’s ability to capture fine-grained semantic distinctions, reinforcing its advantage in retrieval tasks that demand precise understanding.
 7 Conclusion
@@
 We believe that the methodologies and principles of RankCLIP will inspire further research and lead to the development of models with a deeper understanding of the intricate interactions between visual and textual data.
-Limitations
-This work has two main limitations.
-First, due to computational constraints, we were not able to perform more baseline comparisons by pretraining additional existing methods.
-And second, we were unable to scale the pretraining process to include larger datasets, such as those used in OpenCLIP (Ilharco et al., 2021).
 Appendix
-Appendix A CLIP Preliminaries
-CLIP (Radford et al., 2021) has been a prominent method for learning detailed multimodal representations through the alignment of images and texts.
+Appendix A Training Procedures
+A.1 Implementation Details
+For CLIP [43], we use the official implementation released by OpenAI111CLIP repository on GitHub: https://github.com/openai/CLIP..
+And for ALIP [58], we also use the official implementation released by the paper authors222ALIP repository on GitHub: https://github.com/deepglint/ALIP..
+As the proposed RankCLIP essentially shares the same model architecture (separate vision, text encoders, projection layer, and a classification head) as CLIP, we build upon the CLIP code repository for our model construction333RankCLIP repository will be released upon acceptance..
+We set the scaling parameters for cross-modal () and in-modal () ranking consistency to 1/16 and 1/16 respectively throughout all the experiments unless otherwise noted.
+All CLIP, ALIP and RankCLIP models are initialized from scratch without loading any existing weights.
+And the embedding sizes for both modalities all project to 1024 across the three models.
+A.2 Training Parameters
+Following CLIP [43], we adopt the ResNet-50 [21] and transformer architectures [13] for image and text encoding, respectively.
+Training is conducted from scratch over 64 epochs using a single NVIDIA A100 GPU, with a batch size of 512, an initial learning rate of 0.0005 employing cosine scheduling, and 10,000 warm-up steps.
+A.3 Training Time Consumption
+we conducted the experiments using the same hardware specifications.
+The table below shows the time consumption for training our RankCLIP and CLIP models with 50K samples from CC3M using a single NVIDIA A100 GPU.
+Time consumption | Dataset size | epochs | batch_size | model_name | |
+CLIP | 1d 2h 54m 48s | 50K | 64 | 512 | RN50 |
+RankCLIP | 1d 1h 4m 23s | 50K | 64 | 512 | RN50 |
+As shown in the table, the difference in time consumption is negligible.
+Interestingly, our method is slightly faster than CLIP, but we think it may be attributed to hardware optimizations or variance.
+Appendix B CLIP Preliminaries
+CLIP [43] has been a prominent method for learning detailed multimodal representations through the alignment of images and texts.
 Given a set of image-text pairs, where denotes an image and is the corresponding text, the goal is to learn representations that map semantically similar images and texts closer in the embedding space, while dissimilar pairs are distanced apart.
@@
 so that the magnitude information is erased and only direction is preserved.
-To align the image and text representations, a contrastive loss function, typically a variant of the InfoNCE loss (Oord et al., 2018), which optimizes the similarity of the matched pair against unmatched pairs, is utilized, i.e.:
+To align the image and text representations, a contrastive loss function, typically a variant of the InfoNCE loss [39], which optimizes the similarity of the matched pair against unmatched pairs, is utilized, i.e.:
 where the first term contrasts images with the texts, the second term contrasts texts with the images, and denotes a temperature scaling parameter that adjusts the concentration of the distribution.
@@
 This observation motivates the development of RankCLIP, which innovates beyond binary pairwise contrasts to consider holistic listwise consistency within and across modalities.
-Appendix B Training Procedures
-B.1 Implementation Details
-For CLIP (Radford et al., 2021), we use the official implementation released by OpenAI111CLIP repository on GitHub: https://github.com/openai/CLIP..
-And for ALIP (Yang et al., 2023), we also use the official implementation released by the paper authors222ALIP repository on GitHub: https://github.com/deepglint/ALIP..
-As the proposed RankCLIP essentially shares the same model architecture (separate vision, text encoders, projection layer, and a classification head) as CLIP, we build upon the CLIP code repository for our model construction333RankCLIP repository will be released upon acceptance..
-We set the scaling parameters for cross-modal () and in-modal () ranking consistency to 1/16 and 1/16 respectively throughout all the experiments unless otherwise noted.
-All CLIP, ALIP and RankCLIP models are initialized from scratch without loading any existing weights.
-And the embedding sizes for both modalities all project to 1024 across the three models.
-B.2 Training Parameters
-Following CLIP (Radford et al., 2021), we adopt the ResNet-50 (He et al., 2016) and transformer architectures (Devlin et al., 2018) for image and text encoding, respectively.
-Training is conducted from scratch over 64 epochs using a single NVIDIA A100 GPU, with a batch size of 512, an initial learning rate of 0.0005 employing cosine scheduling, and 10,000 warm-up steps.
+B.1 Additional Experiments
+We conduct the linear probing experiment under different training datasize from 3m to 15m as shown in 7.
+Data Size | Method | Model Type |
+CIFAR-10 |
+CIFAR-100 |
+DTD |
+FGVGAircraft |
+Food101 |
+GTSRB |
+OxfordPets |
+SST2 |
+STL10 |
+SVHN |
+CLIP | RN50 | 80.12% | 58.50% | 57.18% | 39.75% | 59.14% | 72.41% | 61.73% | 54.48% | 86.01% | 58.92% | |
+3m | RankCLIP | RN50 | 78.29% | 56.24% | 57.82% | 39.30% | 58.63% | 74.13% | 64.35% | 55.02% | 86.69% | 60.68% |
+CLIP | ViT-B/32 | 77.60% | 56.15% | 43.19% | 22.59% | 39.72% | 62.05% | 40.39% | 50.96% | 78.99% | 50.53% | |
+RankCLIP | ViT-B/32 | 78.42% | 56.64% | 42.39% | 23.43% | 40.19% | 60.63% | 40.56% | 53.32% | 79.60% | 47.72% | |
+CLIP | RN50 | 78.81% | 56.32% | 61.49% | 25.83% | 61.64% | 68.76% | 60.37% | 55.57% | 89.82% | 47.99% | |
+15m | RankCLIP | RN50 | 83.27% | 62.96% | 65.96% | 32.19% | 68.11% | 74.25% | 67.40% | 56.34% | 94.20% | 53.06% |
+CLIP | ViT-B/32 | 82.97% | 62.55% | 49.47% | 24.48% | 52.46% | 63.55% | 50.78% | 52.66% | 87.14% | 46.38% | |
+RankCLIP | ViT-B/32 | 82.79% | 59.89% | 52.50% | 23.94% | 56.44% | 61.58% | 52.98% | 53.60% | 89.01% | 42.16% |